2 trains leave town at the same time and travel in opposite directions. One train travels at
44mph and the other travels at 58mph. How long would it take for the trains to be 561
miles apart?

Answer by Alan3354(69443)

2 trains leave town at the same time and travel in opposite directions. One train travels at 44mph and the other travels at 58mph.
How long would it take for the trains to be 561 miles apart?
----------
The speeds are added when going in opposite directions ---> 102 mi/hr
---
t = d/r = 561/102 hours